If the Demand function is P=10-3q than what will be the total revenue at q=3


1
Expert's answer
2022-01-29T11:05:57-0500

If the Demand function is P = 10 - 3q, then the total revenue at q = 3 is:

"TR = P\u00d7q = (10 - 3\u00d73)\u00d73 = 3."